In moles of acetaldehyde
(CH CHO) calculate the following
a. Number of moles of carbon
b. Number of moles of hydrogen
c. Number of moles of oxygen
d. Number of molecules of acetaldehyde
(Ans.: A. 4 mol, B. 8 mol, C. 2 mol,
D. 12.044 x 1023 molecules )

Grade:12

1 Answers

abhay gupta
3 years ago
Answer:

Therefore, 2 mol of CH3-CHO contains 2×4 = 8 moles of Hydrogen. c) 1 molecule of CH3-CHO contains 1 O-atom. Therefore, 2 mol of CH3-CHO contains 2×1 = 2 moles of Carbon. d) 1 mole of acetaldehyde contains 6.022×10^23 molecules, thus 2 mol will contain 2×6.022×10^23 = 1.2044×10^24 Acetaldehyde molecules

thank you dear !